OpenStudy (eiliasmctalun):

Reina and Sam are multiplying (8^4)^5 (7^3)^9. Reina's Work (8^4)^5 (7^3)^9 = 8^4 + 5^7^3 + ^9 = 8^97^12 Sam's Work (8^4)^5 (7^3)^9 = 8^4⋅5^7^3⋅^9 = 8^207^27 Is either of them correct? Explain your reasoning.

jhonyy9 (jhonyy9):

you missed some signs from there so like result i see the Sam´s work correct but this just than you push,take the product sign between 20 and 7 bc. (8^4)^5 = 8^20 and (7^3)^9 = 7^27 and in this way the right,correct answer will be 8^20 *7^27 hope helped

OpenStudy (niksbhalla14mar):

As per what the question mentioned reflects the equation is (8^4)^5 (7^3)^9 which can be equated using BODMAS(First Solve Brackets) to 4096^5 . 343^9 or 8^20 . 7^21. Here i am assuming that if no sign is mentioned between it is a multiplication sign and so i have used . Hope it helped.